# Multi-Agent Q-Learning Dynamics in Random Networks: Convergence due to Exploration and Sparsity

### v1

@@ Abstract
Beyond specific settings, many multi-agent learning algorithms fail to converge to an equilibrium solution, and instead display complex, non-stationary behaviours such as recurrent or chaotic orbits. In fact, recent literature suggests that such complex behaviours are likely to occur when the number of agents increases. In this paper, we study Q-learning dynamics in network polymatrix games where the network structure is drawn from classical random graph models. In particular, we focus on the Erdős-Rényi model, a well-studied model for social networks, and the Stochastic Block model, which generalizes the above by accounting for community structures within the network. In each setting, we establish sufficient conditions under which the agents’ joint strategies converge to a unique equilibrium. We investigate how this condition depends on the exploration rates, payoff matrices and, crucially, the sparsity of the network. Finally, we validate our theoretical findings through numerical simulations and demonstrate that convergence can be reliably achieved in many-agent systems, provided network sparsity is controlled.
@@ 1 Introduction
The development of algorithms for multi-agent learning has produced a number of successes in recent years, solving challenging problems in load-balancing (Krichene et al., 2014; Southwell et al., 2012; Moghaddam & Kebriaei, 2024), energy management (Paccagnan et al., 2018; Maddouri et al., 2018) and game playing (Moravčík et al., 2017; Brown & Sandholm, 2019; Samvelyan et al., 2019; Perolat et al., 2022). Also thanks to these successes, the game-theoretical foundations of learning in the face of many agents remain a thriving area of research. An important question is whether learning algorithms are guaranteed to converge to an equilibrium solution, often the Nash Equilibrium. However, such guarantees exist only for specialized settings, such as two-player games (Metrick & Polak, 1994; Kianercy & Galstyan, 2012), zero-sum network games (Ewerhart & Valkanova, 2020; Leonardos et al., 2024) or potential games (Krichene et al., 2015; Harris, 1998). Often, such assumptions are restrictive and do not hold in practice. As the development of learning in the presence of many agents progresses, it therefore becomes necessary to understand whether convergence guarantees hold outside these restrictive assumptions.
@@
Unfortunately, it is precisely at this point where results on non-convergence appear. A strong example is Milionis et al. (2023), which introduced a game in which no independent learning dynamics converges to a Nash Equilibrium. Further studies have shown that chaotic dynamics may occur even in two-player finite-action games (Sato et al., 2002; Galla & Farmer, 2013). Crucially, Sanders et al. (2018) found that, as the number of players grows, non-convergent behaviour becomes the norm.
@@
Whilst at first glance these results seem to suggest an insurmountable barrier towards strong guarantees of convergence, a key missing factor is an in-depth analysis of the interactions between agents. In particular, both Sanders et al. (2018) and Hussain et al. (2023) assume that the payoff to any given agent is dependent on all other agents in the environment. Yet, in practice, this assumption rarely holds. Rather, agents often interact through an underlying network that may represent social connections, communication constraints, or spatial proximity. In this work, we address this concern by studying learning in network games (Kearns et al., 2001). In particular, we are motivated by the following research question:
@@
How does network structure affect the convergence of multi-agent learning as the number of agents increases?
@@
Answering this question represents a key step towards guaranteeing the feasibility of learning with many agents, so long as the network structure can be controlled. Indeed, a number of works have examined learning in network games, uncovering the relationship between the network structure and properties of the equilibrium as well as designing algorithms that converge to an equilibrium (Melo, 2018; Parise & Ozdaglar, 2019; Melo, 2021; Shokri & Kebriaei, 2020). However, these algorithms require full knowledge of the agents’ payoff functions or evaluations of its gradient, which may not be available in practice. Our goal instead is to consider a widely applied reinforcement learning algorithm – Q-Learning – which requires only evaluations of the payoff functions, and study how the parameters of the algorithm and the structure of the underlying network can be leveraged to yield convergence guarantees, as well as how such guarantees scale with the number of agents.
@@
We study agents who update their strategies via the Q-Learning dynamics (Sato & Crutchfield, 2003; Tuyls et al., 2006), a continuous-time counterpart to the well-studied Q-Learning algorithm (Watkins & Dayan, 1992; Sutton & Barto, 2018). We focus our study on network polymatrix games (Janovskaja, 1968) in which agents select from a finite set of actions and interact with their neighbours on an underlying network. We study networks which are drawn from random graph models, where the sparsity of the network is parameterised by the distribution from which the network is drawn: the Erdős-Rényi model (Erdös & Rényi, 2006; Newman, 2010) and the Stochastic Block Model (Holland et al., 1983). The Erdős-Rényi model is widely used to study social networks and natural phenomena such as information diffusion Li et al. (2019). The Stochastic Block Model is a natural extension of Erdős-Rényi to model community structures which naturally arise in distributed systems (Yun & Proutière, 2019).
@@
Our main contribution is to provide a precise characterisation of the convergence properties of Q-Learning, depending on the exploration rates of the agents, the payoffs in the game, and the sparsity of the network. Specifically, our results provide sufficient conditions for the Q-learning dynamics to converge to an equilibrium with high probability as the number of agents increases, given that the network is drawn from one of the two above-mentioned models. We show that, in sparsely connected networks, Q-learning converges even with low exploration rates as the number of agents increases, whereas, in agreement with previous results, learning in densely connected networks may not converge. In doing so, we show an explicit relationship between the convergence the Q-Learning Dynamics and the sparsity of the network, thus ensuring the feasibility of learning in many-agent games, so long as the network sparsity is controlled. To the best of our knowledge, ours is the first work to study the asymptotic behaviour of learning dynamics in the context of network polymatrix game with a random graph model, and to derive the relation between the sparsity of the network.
As part of our results, we show that our sufficient conditions also guarantee the uniqueness of the equilibrium. In this way, we show not only that Q-Learning converges, but also that it does so to a unique solution, for all initial conditions. We finally adapt our results to determine a condition that guarantees the uniqueness of the Nash Equilibrium.
Related Work
@@
Our work is focused on independent online learning in network polymatrix games. Network games are well-studied within the setting of zero-sum networks (Cai et al., 2016), which model strictly competitive systems. In such cases, it is known that the continuous-time counterparts of popular algorithms such as Fictitious Play (Ewerhart & Valkanova, 2020) and Q-Learning (Leonardos et al., 2024) converge to an equilibrium. By contrast it was shown in Bailey & Piliouras (2019) and Abernethy et al. (2021) that the Replicator Dynamics (Maynard Smith, 1974), a continuous-time model of the Multiplicative Weights Update algorithm (Arora et al., 2012) does not converge to the NE, but rather cycles at a fixed distance away from it.
@@
Network Games have also been studied to understand the properties, in particular the uniqueness, of the equilibrium (Bramoullé et al., 2014; Parise & Ozdaglar, 2019; Melo, 2021). In many of these cases, the literature appeals to the study of monotone games (Paccagnan et al., 2018) which subsumes zero-sum network games (Akin & Losert, 1984). The formalism of monotone games has been applied to design algorithms that provably converge to Nash Equilibria. However, many of these require that agents have full access to their payoff function (Parise & Ozdaglar, 2019) or its evaluations of its gradient (Mertikopoulos & Zhou, 2019). Monotone games also share strong links with the idea of payoff perturbations (Facchinei & Pang, 2004) in which a strongly convex penalty is imposed to agents’ payoff functions to stabilise learning (Abe et al., 2024; Sokota et al., 2023; Liu et al., 2023).
@@
Our work departs from the above by considering Q-Learning Dynamics, a foundational exploration and exploitation model central to reinforcement learning theory (Albrecht et al., 2024a; Tuyls, 2023). Q-Learning Dynamics is also known to be related to the replicator dynamics (Bloembergen et al., 2015) and to Follow-the-Regularised-Leader (Abe et al., 2022). Outside of specialised classes of games, Q-Learning Dynamics has been shown to exhibit chaotic orbits (Sato et al., 2002; Galla & Farmer, 2013), a phenomenon which becomes more prevalent as the number of players increases (Sanders et al., 2018). Similar to our work, Hussain et al. (2024) consider the convergence of Q-Learning in network polymatrix games. However, their work considered deterministic graphs with specific structures. We instead consider Q-Learning in a stochastic setting, where the stochasticity comes from the random network. In doing so, we derive a direct relationship between the sparsity of edge connections in the network and the convergence of Q-Learning Dynamics.
To achieve our result, we appeal to the framework of random networks, which are often used to model social and community structures. Many works in multi-agent learning, and online optimisations have considered this setting, for example to determine the existence of pure Nash equilibria (Daskalakis et al., 2011), to study the performance of distributed online algorithms (Lei et al., 2020), or to examine the emergence of cooperative behaviours (Dall’Asta et al., 2012). An important step in the study of network games was the introduction of Graphon games in Parise & Ozdaglar (2023), which are the limit of network games as the agents tend to infinity. The authors show that graphons generalise Erdős-Rényi and Stochastic Block Models. However, the restriction to the limit of infinite players is not suitable for our purposes, as it does not allow us to examine how our convergence guarantees are influenced by the number of agents.
2 Preliminaries
We begin in Sec. 2.1 by describing network games. In Sec.  2.2, we describe the equilibrium concepts which are of particular interest in multi-agent learning. We then describe Q-Learning in Sec. 2.3.
2.1 Game Model
@@
In this work, we consider network polymatrix games Janovskaja (1968); Cai et al. (2016), which are defined as tuples $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A}_{k})_{k\in\mathcal{N}},(A^{kl},A^{lk})_{(k,l)\in\mathcal{E}})$, in which $\mathcal{N}=\{1,2,\ldots,N\}$ denotes a set of $N$ agents, indexed by $k$. The interactions between agents is modelled by a set $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$ of edges that defines an undirected network. An alternative formulation of the underlying network is through an adjacency matrix $G\in\mathbb{R}^{N\times N}$ in which $[G]_{kl}=1$ if $(k,l)\in\mathcal{E}$, and $[G]_{kl}=0$ otherwise.
@@
The goal of each agent is to maximise a utility function $u_{k}$. In a network polymatrix game $\mathcal{G}$, each edge is associated with the payoff matrices $(A^{kl},A^{lk})$ so that the payoff for each agent $k$ takes the form
@@
2.2 Solution Concepts
In this work, we are concerned with two widely-studied solution concepts: the Nash Equilibrium (Nash, 1950) and the Quantal Response Equilibrium (McKelvey & Palfrey, 1995). To define these concepts we first define the reward to agent $k\in\mathcal{N}$ for playing action $i\in\mathcal{A}_{k}$ as $r_{ki}(\mathbf{x}_{-k})=\sum_{l:(k,l)\in\mathcal{E}}\sum_{j\in\mathcal{A}_{l}}[A^{kl}]_{ij}x_{lj}$.
Definition 1 (Nash Equilibrium).
A joint strategy $\mathbf{x}^{*}\in\Delta(\mathcal{A})$ is a Nash Equilibrium if, for all $k\in\mathcal{N}$,
$\mathbf{x}_{k}^{*}\in\operatorname*{arg\,max}_{\mathbf{x}_{k}\in\Delta(\mathcal{A}_{k})}u_{k}(\mathbf{x}_{k},\mathbf{x}_{-k}^{*}).$
Definition 2 (Quantal Response Equilibrium (QRE)).
Let $T_{1},\ldots,T_{N}\geq 0$. Then, a joint strategy $\mathbf{x}^{*}\in\Delta(\mathcal{A})$ is a Quantal Response Equilibrium if, for all $k\in\mathcal{N}$,
$\mathbf{x}_{k}^{*}=\frac{\exp\left(r_{ki}(\mathbf{x}_{-k})/T_{k}\right)}{\sum_{j\in\mathcal{A}_{k}}\exp\left(r_{kj}(\mathbf{x}_{-k})/T_{k}\right)}.$
The QRE is a natural extension of the Nash Equilibrium that allows for agents to play suboptimal actions with non-zero probability. This is crucial in online learning, where agents must explore their strategy space. The parameter $T_{k}$ is therefore known as the exploration rate of agent $k$. Notice that, by taking the limit $T_{k}\rightarrow 0$ for all $k$, the QRE converges to the Nash Equilibrium (McKelvey & Palfrey, 1995). On the other hand, as $T_{k}\rightarrow\infty$ for all $k$, the QRE converges to the uniform distribution, in which payoffs are ignored and agents play their actions at random. For this reason, it is preferred that the agents learn QRE at lower values of $T_{k}$, so that the learned solution approaches the Nash Equilibrium (Hussain et al., 2024).
2.3 Q-Learning Dynamics
We now describe the Q-Learning algorithm (Watkins & Dayan, 1992) which aims to learn an optimal action-value function $Q:\mathcal{A}\rightarrow\mathbb{R}$ that captures the expected reward of taking a given action. We consider the multi-agent extension of Q-Learning (Schwartz, 2014; Albrecht et al., 2024b) in which each agent $k$ maintains an individual Q-value estimate $Q_{k}:\mathcal{A}_{k}\rightarrow\mathbb{R}$. These are updated at each round $t$ via the update
$Q_{ki}(t+1)=(1-\alpha_{k})Q_{ki}(t)+\alpha_{k}r_{ki}(\mathbf{x}_{-k}(t))$
@@ 2.3 Q-Learning Dynamics
In effect, $Q_{ki}$ gives a discounted history of the rewards received when action $i$ is played, with $1-\alpha_{k}$ as the discount factor. It should also be noted that Q-values are updated by the rewards $r_{ki}(\mathbf{x}_{-k})$ that depend on the strategies of all other agents at time $t$. So, the reward for a single action can vary between rounds. It is this non-stationarity that often leads to chaotic dynamics in multi-agent learning (Sato et al., 2002).
Given the Q-values, each agent plays their actions according to the Boltzmann distribution (Sutton & Barto, 2018) given by
$x_{ki}(t+1)=\frac{\exp\left(Q_{ki}(t+1)/T_{k}\right)}{\sum_{j\in\mathcal{A}_{k}}\exp\left(Q_{kj}(t+1)/T_{k}\right)}$ (1)
@@
Tuyls et al. (2006) and Sato & Crutchfield (2003) showed that a continuous-time approximation of the Q-Learning algorithm is given by a variation of the replicator dynamics (Maynard Smith, 1974; Hofbauer & Sigmund, 2003) that is called the Q-Learning Dynamics
$\frac{\dot{x}_{ki}}{x_{ki}}=r_{ki}\left(\mathbf{x}_{-k}\right)-\langle\mathbf{x}_{k},r_{k}(\mathbf{x})\rangle+T_{k}\sum_{j\in\mathcal{A}_{k}}x_{kj}\ln\frac{x_{kj}}{x_{ki}},$ (QLD)
[…]
@@
Given a square matrix $A\in\mathbb{R}^{N\times N}$, we denote its spectral norm as $\lVert A\rVert_{2}=\sup_{x\in\mathbb{R}^{N}\,:\,\lVert\mathbf{x}\rVert_{2}=1}\lVert Ax\rVert_{2}$. If $A$ is symmetric, we denote its spectral radius as $\rho(G)=\max_{i=1,\ldots,N}|\lambda_{i}|$, where $\lambda_{1},\ldots,\lambda_{N}$ denote the eigenvalues of $A$.

### v2

# Convergence and Connectivity: Dynamics of Multi-Agent Q-Learning in Random Networks

@@ Abstract
Beyond specific settings, many multi-agent learning algorithms fail to converge to an equilibrium solution, instead displaying complex, non-stationary behaviours such as recurrent or chaotic orbits. In fact, recent literature suggests that such complex behaviours are likely to occur when the number of agents increases. In this paper, we study Q-learning dynamics in network polymatrix normal-form games where the network structure is drawn from classical random graph models. In particular, we focus on the Erdős-Rényi model, which is used to analyze connectivity in distributed systems, and the Stochastic Block model, which generalizes the above by accounting for community structures that naturally arise in multi-agent systems. In each setting, we establish sufficient conditions under which the agents’ joint strategies converge to a unique equilibrium. We investigate how this condition depends on the exploration rates, payoff matrices and, crucially, the probabilities of interaction between network agents. We validate our theoretical findings through numerical simulations and demonstrate that convergence can be reliably achieved in many-agent systems, provided interactions in the network are controlled.
@@ 1 Introduction
The development of algorithms for multi-agent learning has produced a number of successes in recent years, solving challenging problems in load-balancing (Krichene et al., 2014; Southwell et al., 2012), energy management (Maddouri et al., 2018) and game playing (Moravčík et al., 2017; Brown and Sandholm, 2019; Samvelyan et al., 2019; Perolat et al., 2022). Also thanks to these successes, the game-theoretic foundations of learning in the face of many agents remain a thriving area of research. As the number of agents grows in these systems, it is critical to understand if their learning algorithms are guaranteed to converge to an equilibrium solution, such as a Nash Equilibrium.
@@
Unfortunately, previous work suggests that non-convergent behaviour becomes the norm as the number of players increases. A strong example is (Milionis et al., 2023), which introduced a game in which no independent learning dynamics converges to a Nash Equilibrium. Further studies have shown that chaotic dynamics may occur even in two-player finite-action games (Sato et al., 2002; Galla and Farmer, 2013). Crucially, Sanders et al. (2018) found that, as the number of players grows, non-convergent behaviour becomes the norm.
@@
Whilst at first glance these results suggest an insurmountable barrier towards strong convergence guarantees, a key missing factor is an in-depth analysis of the interactions between agents. In particular, both Sanders et al. (2018) and Hussain et al. (2023) assume that the payoff to any given agent is dependent on all other agents in the environment. This assumption rarely holds in practice. Rather, agents often interact through an underlying network that may represent communication constraints or spatial proximity. A practitioner or system designer has a certain degree of control over the network structure, e.g., in a robotic swarm, they can influence the network by adjusting the communication range of the robots, while in a sensor network, this can be done by controlling the density of the deployed sensors. This leads us to study the following research question:
@@
How does network structure affect the convergence of learning as the number of agents increases?
@@
Answering this question represents a key step towards guaranteeing the feasibility of learning with many agents, so long as the network structure can be controlled. Indeed, a number of works have examined learning in network games, uncovering the relationship between the network structure and properties of the equilibrium as well as designing algorithms that converge to an equilibrium (Melo, 2018; Parise and Ozdaglar, 2019; Melo, 2021; Shokri and Kebriaei, 2020). However, many of these algorithms have unrealistic requirements, such as full knowledge of the agents’ payoff functions or their gradients. Our goal instead is to consider a widely applied reinforcement learning algorithm – Q-Learning – which requires only evaluations of the payoff functions, and study how the parameters of the algorithm and the structure of the underlying network can be leveraged to yield convergence guarantees, as well as how such guarantees scale with the number of agents.
@@
We study agents who update their strategies via the Q-Learning dynamics (Sato and Crutchfield, 2003), a continuous-time counterpart to the well-studied Q-Learning algorithm (Watkins and Dayan, 1992; Sutton and Barto, 2018). We focus our study on network polymatrix games (Janovskaja, 1968) in which agents select from a finite set of actions and interact with their neighbours on an underlying network. Network polymatrix games provide a setting in which the topology of interaction directly shapes both payoff structure and learning dynamics, allowing us to isolate and quantify the role of network structure in the convergence of multi-agent online learning. We study networks which are drawn from random graph models: the Erdős-Rényi model (Erdös and Rényi, 2006) and the Stochastic Block model (Holland et al., 1983). The Erdős-Rényi model is widely used to study communication in distributed systems (Lei et al., 2020; Prakash et al., 2020). The Stochastic Block model is a natural extension of Erdős-Rényi to model community structures which naturally arise in distributed systems (Yun and Proutière, 2019).
To prove our main results (Theorems 1 and 2), we restrict to network polymatrix games in which each edge of the network is assigned the same bimatrix game. This means that agents are engaged in the same interaction with multiple opponents. This assumption is commonly used to study the emergence of cooperation (Zhang et al., 2014; Mukhopadhyay and Chakraborty, 2020) and congestion (Szabó and Szolnoki, 2015) in multi-agent systems. Indeed, the assumption of shared payoffs has led to a number of successes in studying large scale systems with many agents (Perrin et al., 2020; Wu et al., 2024; Pérolat et al., 2022; Parise and Ozdaglar, 2023; Hu et al., 2019). We adopt this assumption, but also empirically validate our claims outside of this framework through the Conflict Network game (Ewerhart and Valkanova, 2020), in which payoff matrices vary across edges.
@@
Our main contribution is to characterise the convergence properties of Q-Learning, depending on the exploration rates of the agents, the game payoffs, and the expected degree of the network. Specifically, our results provide sufficient conditions for the Q-learning dynamics to converge to an equilibrium with high probability as the number of agents $N$ increases, provided that the network is drawn from one of the two above-mentioned models. We demonstrate that in low-degree networks, Q-learning converges even with low exploration rates and with a large number of agents. In fact, our experiments show convergence even with $200$ agents in low-degree networks, whereas Sanders et al. (2018) reported failure to converge with low exploration rates for as few as five agents.
Our work establishes an explicit relationship between the convergence of Q-Learning Dynamics and the expected node degree in the network, thus ensuring the feasibility of learning in many-agent games, so long as the expected node degree is controlled. Our results further ensure the uniqueness of the equilibrium, meaning that Q-learning converges to a single solution regardless of the initial conditions. To the best of our knowledge, this is the first work to study the asymptotic behaviour of learning dynamics in the context of network polymatrix games with random graph models, and to derive the relation between convergence and the expected node degree in the network. For reasons of space, the proofs of all results appear in the Appendix.
1.1 Related Work
@@
Our work focuses on independent online learning in network polymatrix games. Network games are well-studied in the setting of zero-sum games (Cai et al., 2016), which model strictly competitive systems. In such cases, it is known that the continuous-time counterparts of popular algorithms such as Fictitious Play (Ewerhart and Valkanova, 2020) and Q-Learning (Leonardos et al., 2024) converge to an equilibrium. By contrast, Bailey and Piliouras (2019) showed that the Replicator Dynamics (Maynard Smith, 1974), a continuous-time model of the Multiplicative Weights Update algorithm (Arora et al., 2012), does not converge to the Nash Equilibrium. Similarly, Shapley (1964) showed the non-convergence of Fictitious Play in a two-person non-zero-sum game. Indeed, Hart and Mas-Colell (2003) showed that no learning dynamic that is uncoupled, in that the strategy of a player is independent of the payoff functions of other players, can converge to a Nash Equilibrium in all normal-form games. Milionis et al. (2023) extended this result to show the impossibility of convergence of a wide class of learning dynamics to approximate Nash-Equilibria in normal-form games.
@@
Network games have also been studied to understand the properties, in particular the uniqueness, of the equilibrium (Bramoullé et al., 2014; Parise and Ozdaglar, 2019; Melo, 2021). In many cases, the literature appeals to the study of monotone games (Paccagnan et al., 2018) which subsumes zero-sum network games (Akin and Losert, 1984). The formalism of monotone games has been applied to design algorithms that provably converge to Nash Equilibria. However, many of these algorithms require that agents have full access to their payoff function (Parise and Ozdaglar, 2019) or its gradient (Mertikopoulos and Zhou, 2019). Monotone games also share strong links with the idea of payoff perturbations (Facchinei and Pang, 2004) in which a strongly convex penalty is imposed to agents’ payoff functions to stabilise learning (Abe et al., 2024; Sokota et al., 2023; Liu et al., 2023).
@@
Our work departs from the above by considering Q-Learning Dynamics, a foundational exploration-exploitation model central to reinforcement learning (Albrecht et al., 2024a; Tuyls, 2023). Q-Learning Dynamics is also related to the replicator dynamics (Bloembergen et al., 2015) and to Follow-the-Regularised-Leader (Abe et al., 2022). Outside of specialised classes of games, Q-Learning Dynamics has been shown to exhibit chaotic orbits (Sato et al., 2002), a phenomenon which becomes more prevalent as the number of players increases (Sanders et al., 2018). Similar to our work, Hussain et al. (2024) study the convergence of Q-Learning in network polymatrix games. However, their work considered deterministic graphs and, as such, derived sufficient conditions for convergence in terms of norms of the network adjacency matrix. We instead consider Q-Learning in a stochastic setting, where stochasticity arises from the random network. In doing so, we derive a direct relationship between the expected node degree and the convergence of Q-Learning Dynamics. Our random-network analysis also allows for the study of certain network classes, e.g., the Stochastic Block Model, which are well-studied models of distributed systems that contain sub-communities. With the tools of random networks, we make inroads into understanding how inter-, or intra- community edge connections affect the resulting convergence of the learning algorithm.
Other works in multi-agent learning and online optimisation have considered the random network setting, for example to determine the existence of pure Nash Equilibria (Daskalakis et al., 2011), to study the performance of distributed online algorithms (Lei et al., 2020), or to examine the emergence of cooperative behaviors (Dall’Asta et al., 2012). A parallel step in the study of network games was the introduction of graphon games in (Parise and Ozdaglar, 2023; Carmona et al., 2022). The authors consider network games in the limit of uncountably infinite agents and generalise the Erdős-Rényi and Stochastic Block models. This also extends mean-field games (Laurière et al., 2022; Hu et al., 2019) by introducing heterogeneity amongst players through their edge connections. Subsequent works on graphon games largely focus on the analysis of equilibria (Caines and Huang, 2021; Aurell et al., 2022) or design learning algorithms that converge in time-average to an equilibrium (Cui and Koeppl, 2022; Zhang et al., 2023, 2024). By contrast, our goal is to understand the last-iterate behaviour of Q-Learning Dynamics and establish probabilistic bounds to guarantee convergence in games with finitely many players. In particular, we study how the probability of edge connections in the network curtails non-convergent dynamics as the number of players increases. Thus, the restriction to uncountably infinite players is not suitable for our purposes, although the analysis of non-convergent learning algorithms in graphon games is an interesting direction for future work.
2 Background
Game Model.
@@ 2 Background
We consider network polymatrix games (Janovskaja, 1968; Cai et al., 2016), which are defined as tuples $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A}_{k})_{k\in\mathcal{N}},(A^{kl},A^{lk})_{(k,l)\in\mathcal{E}})$, where $\mathcal{N}=\{1,2,\ldots,N\}$ denotes a set of $N$ agents, indexed by $k$. The interactions between agents are modelled by a set $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$ of edges that defines an undirected network. An alternative formulation of the underlying network is through an adjacency matrix $G\in\mathbb{R}^{N\times N}$ in which $G]_{kl}=1$ if $(k,l)\in\mathcal{E}$, and $G]_{kl}=0$ otherwise. The degree of an agent $k\in\mathcal{N}$ is the number of edges in $\mathcal{E}$ that include $k$.
@@
The goal of each agent is to maximise a utility function $u_{k}$. In a network polymatrix game $\mathcal{G}$, each edge is associated with the payoff matrices $(A^{kl},A^{lk})$, i.e., the payoff for each agent $k\in\mathcal{N}$ is
@@
Solution Concepts.
We focus on two widely-studied solution concepts: the Nash Equilibrium (Nash, 1950) and Quantal Response Equilibrium (McKelvey and Palfrey, 1995). To define these concepts we first introduce the reward to agent $k\in\mathcal{N}$ for playing action $i\in\mathcal{A}_{k}$ as $r_{ki}(\mathbf{x}_{-k})=\sum_{l:(k,l)\in\mathcal{E}}\sum_{j\in\mathcal{A}_{l}}[A^{kl}]_{ij}x_{lj}$.
Definition 1 (NE).
A joint strategy $\mathbf{x}^{*}\in\Delta(\mathcal{A})$ is a Nash Equilibrium if
$\mathbf{x}_{k}^{*}\in\operatorname*{arg\,max}_{\mathbf{x}_{k}\in\Delta(\mathcal{A}_{k})}u_{k}(\mathbf{x}_{k},\mathbf{x}_{-k}^{*})\text{ for all }k\in\mathcal{N}.$
Definition 2 (QRE).
Let $T_{1},\ldots,T_{N}>0$. Then, a joint strategy $\mathbf{x}^{*}\in\Delta(\mathcal{A})$ is a Quantal Response Equilibrium if
$\mathbf{x}_{k}^{*}=\frac{\exp\left(r_{ki}(\mathbf{x}_{-k}^{*})/T_{k}\right)}{\sum_{j\in\mathcal{A}_{k}}\exp\left(r_{kj}(\mathbf{x}_{-k}^{*})/T_{k}\right)}\text{ for all }k\in\mathcal{N}.$
The QRE is a natural extension of the Nash Equilibrium that allows agents to play suboptimal actions with non-zero probability. This is crucial in online learning, where agents must explore their strategy space. The parameter $T_{k}$ is therefore known as the exploration rate of agent $k$. Notice that, by taking the limit $T_{k}\rightarrow 0$ for all $k$, the QRE converges to the Nash Equilibrium (McKelvey and Palfrey, 1995).
Q-Learning Dynamics.
We now describe the Q-Learning algorithm (Watkins and Dayan, 1992; Sutton and Barto, 2018) which aims to learn an optimal action-value function $Q:\mathcal{A}\rightarrow\mathbb{R}$ that captures the expected reward of taking a given action. To do this, an agent must explore their action space by playing possibly suboptimal actions in an attempt to discover optimal strategies. Simultaneously, the agent plays optimal actions based on their current Q-estimate. As such, the Q-learning algorithm is a foundational model to describe the behaviour of agents that must balance exploration and exploitation. We consider the multi-agent extension of Q-Learning (Schwartz, 2014; Albrecht et al., 2024b) in which each agent $k$ maintains an individual Q-value estimate $Q_{ki}\in\mathbb{R}$ for each of their actions $i\in\mathcal{A}_{k}$. These are updated at each round $t$ via the update
$Q_{ki}(t+1)=(1-\alpha_{k})Q_{ki}(t)+\alpha_{k}r_{ki}(\mathbf{x}_{-k}(t)),$ (1)
@@
In effect, $Q_{ki}$ gives a discounted history of the rewards received when action $i$ is played, with $1-\alpha_{k}$ as the discount factor. Note that Q-values are updated by the rewards $r_{ki}(\mathbf{x}_{-k})$ that depend on the strategies of all other agents at time $t$. So, the reward for a single action can vary between rounds. This non-stationarity often leads to chaotic dynamics in multi-agent learning.
Given the Q-values, each agent plays their actions according to the Boltzmann distribution:
$x_{ki}(t+1)=\frac{\exp\left(Q_{ki}(t+1)/T_{k}\right)}{\sum_{j\in\mathcal{A}_{k}}\exp\left(Q_{kj}(t+1)/T_{k}\right)},$ (2)
@@
Tuyls et al. (2006) and Sato and Crutchfield (2003) showed that a continuous-time approximation of the Q-Learning algorithm is given by a variation of the replicator dynamics (Maynard Smith, 1974; Hofbauer and Sigmund, 2003) that is called the Q-Learning Dynamics
$\frac{\dot{x}_{ki}}{x_{ki}}=r_{ki}\left(\mathbf{x}_{-k}\right)-\langle\mathbf{x}_{k},r_{k}(\mathbf{x})\rangle+T_{k}\sum_{j\in\mathcal{A}_{k}}x_{kj}\ln\frac{x_{kj}}{x_{ki}}.$ (QLD)
where $\langle\cdot,\cdot\rangle$ denotes the scalar product. Leonardos et al. (2024) proved that the fixed points of (QLD) coincide with the QRE of the game.
@@
Given a square matrix $A\in\mathbb{R}^{N\times N}$, we denote its spectral norm as $\lVert A\rVert_{2}=\sup_{x\in\mathbb{R}^{N}\,:\,\lVert\mathbf{x}\rVert_{2}=1}\lVert Ax\rVert_{2}$. If $A$ is symmetric, all its eigenvalues $\lambda_{1},\ldots,\lambda_{N}$ are real, and the spectral norm agrees with the spectral radius, which is defined as $\rho(G)=\max\{\lambda_{1},\ldots,\lambda_{N}\}.$